OpenStudy (anonymous):

The height, h, in meters above the ground, of a projectile at any time, t, in seconds, after launch is defined by the function h = -4t^2+ 48t+ 3.l Find the maximum height reached by the projectile.

OpenStudy (campbell_st):

do you know calculus...?

OpenStudy (anonymous):

agh i need help with some explanation, as I'm doing this course online and the tutorials aren't helping much :/

OpenStudy (anonymous):

Then watch this prof : http://www.youtube.com/watch?v=jbIQW0gkgxo he will give the concept very well

OpenStudy (mathmate):

Completing the square will give you the maximum height (at the vertex): \[h(t)=-4t^2+48t+3=-4(t-6)^2+147=-4(t-h)^2+k\] where (h,k) is the vertex, or where the projectile reaches its maximum. => (h,k)=(6,147) The projectile reaches 147 m, same result if calculus was used. If completing the square appears difficult, you can always take the short cut in that the value of t at the vertex = -b/2a = -48/(2*-4) = 6, and find h(6) accordingly.

OpenStudy (mathmate):



OpenStudy (anonymous):

ahhh thanks for your help! :D

OpenStudy (anonymous):

i really appreciate it xD